OpenStudy (anonymous):

In using the Egyptian doubling procedure to find the product 10 x 104 which lines below should be added to find the product. Line 1 1 104 Line 2 2 208 Line 3 4 312 Line 4 8 416

OpenStudy (amistre64):

that doesnt look correct to me one side is doubled while the other is halved is what i recall

OpenStudy (amistre64):

\[24*7\] \[\frac 22*24*7\] \[\frac {24}2*2(7)\] \[12*14\] forget which gets which .... but it works out to the odd man out type construct

OpenStudy (amistre64):

.... and what product are you trying to determine?

OpenStudy (anonymous):

10x104

OpenStudy (amistre64):

lets see if i recall the process :) 10 104 20 52 40 26 80 13 160 6 +1 320 3 640 1 +1 160 640 ----- 800 , well that wasnt it :/

OpenStudy (amistre64):

can you correct the process any?

OpenStudy (amistre64):

10 104 20 52 40 26 80 13 160 6 + 160 1/2 320 3 640 1 + 640 1/2 640 + 640/2 + 160/2 is the results so every odd value, 1,3, and 13 get added

OpenStudy (amistre64):

if we change the odds to a +1, then its the rows with 1s next to them 10 104 20 52 40 26 80 12 +1 160 6 320 2 +1 640 1

OpenStudy (anonymous):

oops i mean 10

OpenStudy (anonymous):

10 = 2 + 8 i.e. 2,8 are the powers of 2 that add to 10

OpenStudy (anonymous):

the problem is that the lines are wrong!

OpenStudy (anonymous):

I am not sure how you do it

OpenStudy (anonymous):

1 104 double to get 2 208 double again 4 416 and again 8 832

OpenStudy (anonymous):

In using the Egyptian doubling procedure to find the product which lines below should be added to find the product. Line 1 1 104 Line 2 2 208 Line 3 4 312 Line 4 8 416 A) lines 2 and 4 B) lines 1, 2, 3, and 4 C) lines 3 and 4 D) line 4 E) lines 1 and 4

OpenStudy (anonymous):

now \(10=2+8\) so add lines 2 and 8

OpenStudy (anonymous):

that was the question and the answers below ok

OpenStudy (anonymous):

well then answers are wrong, as it the procedure you are supposed to double each time it is pretty clear that \(10\times 104=1040\)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

the largest power of 2 in 10 is 8 then the next largest is 2 and you are done because \(8+2=10\)

OpenStudy (anonymous):

thats i thought

OpenStudy (anonymous):

and if you look at the procedure i wrote above, you add line 8 to line 2 and get \(832+208=1040\)

OpenStudy (anonymous):

it is not clear at all where the 312 comes from in the question, because that is not twice 208 i think there is just a mistake here somewhere in the question

OpenStudy (anonymous):

I think there is too. Let me email the teacher and ask her

OpenStudy (anonymous):

I just emailed the teacher so I guess i will wait on that question. Thank you.